Find the equation of a circle with center at (-1,-3) passing through the intersection point of the two lines: 4x+y-4=0 and x-y-6=0


1
Expert's answer
2021-10-28T11:50:56-0400

Equations of lines are:

"4x+y-4=0\ldots(1)\\ x-y-6=0\ldots(2)"

From "(1)" and "(2)", we get:

"y=-4,x=2"

"\therefore" Point of intersection is "(2,-4)."

Now, center of circle is "(-1,-3)."

"\therefore" Radius is "\sqrt{(2+1)^2+(-4+3)^2}=\sqrt{10}"

Equation of circle is:

"(x+1)^2+(y+3)^2=10\\
\Rightarrow x^2+y^2+2x+6y=0"